OpenStudy (anonymous):

Find the value of x to the nearest tenth. (give me a minute to draw the triangle)

OpenStudy (anonymous):



OpenStudy (anonymous):

3x-10 is the midsegment, so 3x-10 = half of 20

OpenStudy (anonymous):

@dlipson1 so is it 6?

OpenStudy (anonymous):

Means you would write 3x-10=10 right? then solve for x

OpenStudy (anonymous):

not 6, but 3x-10=10 is correct.

OpenStudy (anonymous):

oh wait my bad it would be 6.6?